A meeting of envoys was attended by 4 Koreans and 4 Filipinos. If three envoys were selected at random one after the other, determine the values of the random variable K representing the number of Koreans.

1
Expert's answer
2022-02-23T14:00:28-0500

Possible outcomes: {KKK, KKF, KFK, FKK, KFF, FKF, FFK, FFF}

"P(K=0)=\frac{1}{8}"

"P(K=1)=\frac{3}{8}"

"P(K=2)=\frac{3}{8}"

"P(K=3)=\frac{1}{8}"